OpenStudy (anonymous):

Which conjunction or disjunction equivalent to the open sentence 4 –3|n + 2| ≥ 1? a. n + 2 ≥ 1 and n + 2 ≤ –1 b. n + 2 ≤ 1 and n + 2 ≥ –1 c. n + 2 ≤ 1 and n − 2 ≥ −1 d.n + 2 ≥ 1 or n – 2 ≤ 1 please help >.<

OpenStudy (anonymous):

i tried, but I'm like so confused.

OpenStudy (misty1212):

suppose you changed it to \[4-3X=1\] a totally different, but similar question could you solve for \(X\) ?

OpenStudy (anonymous):

yeah, that makes sense c:

OpenStudy (misty1212):

i hope so, but the question is, can you solve that for \(X\) if so, we can go from there

OpenStudy (anonymous):

it would be x = 1 , right?

OpenStudy (misty1212):

yes it would!

OpenStudy (misty1212):

now that is the question, but the method you used to get \(X=1\) is the same method you would use to get at the answer but since this is an inequality we have to be more careful

OpenStudy (misty1212):

\[4-3|n+2|\geq 1\\ -3|n+2|\geq -3\]and here comes the "be careful" part when you divide by \(-3\) you have to change the inequality

OpenStudy (misty1212):

\[-3|n+2|\geq -3\\ |n+2|\leq 1\]

OpenStudy (misty1212):

now you can solve \[-1\leq n+2\leq 1\] in one easy step by subtracting 2 all the way across

OpenStudy (anonymous):

okay, thank you again!

OpenStudy (misty1212):

you are welcome again!

OpenStudy (misty1212):

i see that they do not (for some unknown reason) want to to solve this, just two write it as a conjunction \[n+2\leq 1\text { and } n +2\geq -1\]not sure why you are not asked to solve